OpenStudy (anonymous):

Perform all operations and express your answer in a + bi form. 7 over 9 + i

jimthompson5910 (jim_thompson5910):

Multiply numerator and denominator by the complex conjugate of 9+i (which is 9-i) to get \[\frac{7(9-i)}{(9+i)(9-i)}\] \[\frac{63-7i}{82}\] So \[\frac{7(9-i)}{(9+i)(9-i)}=\frac{63-7i}{82}\]